40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Apr 28 11:19:24 2022, shruti, Update, PSOMA, Switch current driver to custom low noise 
    Reply  Thu Apr 28 14:44:28 2022, aaron, Update, PSOMA, Switch current driver to custom low noise fitted_full.pdfCalibratedNoiseBudget.pdf
       Reply  Fri Apr 29 11:21:51 2022, shruti, Update, PSOMA, Switch current driver to custom low noise CalibratedNoiseBudget.pdflaser_noise_comparison.pdf
Message ID: 2886     Entry time: Fri Apr 29 11:21:51 2022     In reply to: 2885
Author: shruti 
Type: Update 
Category: PSOMA 
Subject: Switch current driver to custom low noise 

Using the delay line setup with a digital PID slow control on the Moku:Pro in multi-instrument mode, I first calibrated the delay line by offset locking and measuring the beat frequency with small offsets, and then measured the noise spectra at the delay line with no offset. The script I used to analysze the results is here.

Then I ran the noise calibration script and plotted this instead of the previously measured data [Attachment 1].

Attachment 2 contains a comparison of the old and new laser noises measured with the North-South beat. The north laser is still powered by the ITC510 which probably is why the forest of peaks still exist.

 

Attachment 1: CalibratedNoiseBudget.pdf  196 kB  Uploaded Fri Apr 29 12:22:57 2022  | Hide | Hide all
CalibratedNoiseBudget.pdf
Attachment 2: laser_noise_comparison.pdf  60 kB  Uploaded Fri Apr 29 12:53:49 2022  | Hide | Hide all
laser_noise_comparison.pdf
ELOG V3.1.3-